# SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation

### v1

@@ 3.3 Post-processing for RAG
We first extract the sub-images of these elements by the coordinates of SCAN’s predicted bounding box result. For textual RAG, we use VLMs to convert these sub-images into text, then combine them according to their original reading order relationships to obtain the page-level result for retrieval. We can use general rule-based reading order methods; in our experiments, we simply sorted the reading order based on the coordinates of the upper-left corners of various elements.
For visual RAG, since we aim to use semantic boxes as new retrieval object units to replace original pages with larger information content, we have three options and ablation in Section 4.4 test the performance of them.
1. all-box: Directly using the blocks segmented after semantic layout analysis as new objects for image retrieval. Note that titles, headers, etc., are treated as independent images.
2. semantic_box-only: Using only the content of semantic_boxes as new objects for image retrieval.
3. concatenation: Using a concatenation method to combine global information with semantic box content. Specifically, we append the sub-images of (title, header, footer, date, and author) to each semantic_box sub-image.
3.4 Cost Discussion for VLM text conversion
When applying SCAN for textual RAG, we replace a single high-resolution document page with small sub-images, but the VLM inference cost is almost unchanged. This is because current mainstream VLMs consider number of tokens based on image pixels. For example, the Qwen-2.5-VL model corresponds 28 $\times$ 28 pixels to a single visual token. SCAN divides images into sub-images, but the total number of pixels remains essentially unchanged, or decreased since SCAN skips some white areas. Therefore, the number of input visual tokens remains roughly the same. Although multiple VLM calls will increase the total text tokens due to repeated text conversion prompts, this increase is negligible since visual tokens dominate the computational cost.
During inference, because of the KV cache mechanism, time efficiency generally increases linearly with the length of output tokens. Therefore, for an expect text conversion result, the cost of SCAN’s multiple inferences with sub-images should be comparable to a single inference with the entire page image. Moreover, SCAN’s segmentation embodies a divide-and-conquer approach that can be parallelized, fully utilizing GPU capabilities when resources permit. Our experiments in Section 4.5 empirically validate this analysis.
@@ 4 Experiments
4.1 Datasets and Setting
We use three datasets to evaluate the RAG performance. One English dataset OHR-Bench Zhang et al. (2025) and two Japanese datasets, Allganize Allganize.ai (2024) and our in-house BizMMRAG. Every dataset is for both textual and visual RAG. Table 2 gives a detailed summary of these datasets.
Dataset Domain Lang. #Docs #Pages #QA QA Type
OHR-Bench Zhang et al. (2025) Academic, Administration, Finance, En 1,261 8,561 8,498 Text, Table, Formula,
Law, Manual, Newspaper, Textbook Chart, Reading Order
Allganize Allganize.ai (2024) Finance, Information Technology, Manufacturing, Ja 64 2,167 300 Text, Table, Chart
Public Sector, Retail/Distribution
BizMMRAG Finance, Government, Medical, Ja 42 3,924 146 Text, Table, Chart
Consulting sectors, Wikipedia
Table 2: Statistics of our evaluation datasets.
@@ 4.1 Datasets and Setting
SCAN-GOT-580M (ours) 60.4 44.1 42.6 28.5 4.5 45.8 ($+$0.4) 45.4 47.4 27.3 20.3 23.2 36.3 ($+$8.5) 38.4 25.2 20.3 15.2 17.4 28.1 ($+$3.5)
SCAN-Nougat-350M (ours) 61.7 27.9 37.1 10.8 4.3 39.5 (--1.4) 41.9 22.7 21.9 6.4 4.1 27.3 ($+$1.8) 37.3 17.9 17.5 6.4 5.4 23.5 ($+$9.0)
@@
DocLayout-YOLO-Qwen2.5-VL-72B 63.5 12.4 36.0 11.7 5.4 36.3 (--23.3) 41.2 10.2 19.0 7.2 18.2 24.4 (--13.1) 38.4 10.0 16.3 7.7 12.5 22.4 (--8.7)
Beehive-Qwen2.5-VL-72B 73.3 16.0 42.3 16.3 6.1 42.6 (--17.0) 46.6 11.3 21.3 10.2 27.2 28.2 (--9.3) 43 10.9 19.5 8.6 16.6 25.3 (--5.8)
SCAN-Qwen2.5-VL-72B (ours) 75.2 56.8 56.6 36.8 6.2 58.3 (--1.3) 47.4 41.7 30.2 23.9 28.3 39.5 ($+$2.0) 43.3 31.3 25.2 17.9 17.5 32.8 ($+$1.7)
Table 3: Textual RAG results on OHR-Bench. Comparison of various OCR methods across different evaluation metrics. TXT, TAB, FOR, CHA, RO, and ALL represent text, table, formula, chart, reading order, and their average, respectively. The RO category includes questions that require identifying the correct reading order to associate information from separate paragraphs.
When evaluating OHR-Bench for textual RAG, we follow the same setting as in the paper Zhang et al. (2025). We use BGE-m3 Chen et al. (2024a) and BM25 as retrieval models, and meta-llama/Llama-3.1-8B-Instruct and Qwen/Qwen2-7B-Instruct as answer models to generate answers according to the retrieved top-2 results. We evaluate three metrics: (a) retrieval, which calculates LCS (Longest Common Subsequence) to measure evidence inclusion in retrieved content; (b) generation, which measures the F1-score of QA when provided with the ground truth page; and (c) overall, which calculates the F1-score of QA for the end-to-end RAG pipeline. The F1-score is calculated using precision and recall of common tokens between the generated result and the ground truth. The final scores will be the average score across the combination of two retrieval models and two answer models. On the other hand, we also apply OHR-Bench to visual RAG, which uses ColQwen2-v1.0 Faysse et al. (2024) as an image retrieval model with top-5 retrieval, and use Qwen/Qwen2.5-VL-7B as an answer model.
For the Japanese datasets BizMMRAG and Allganize, we employ intfloat/multilingual-e5-large as a textual retrieval model, and ColQwen2-v1.0 as an image retrieval model. For answer generation, we employ GPT-4o (gpt-4o-2024-08-06) in textual RAG and Qwen2.5-VL-7B in visual RAG. For each query, the top-5 retrieved results are used. For evaluation, we adopt the LLM-as-a-judge framework Gu et al. (2025), using GPT-4o with a temperature of 0.3 to assign an integer score from 1 to 5 to each generated answer. Answers receiving a score greater than 4 are considered correct (assigned a value of 1), while others are considered incorrect (assigned a value of 0). Final accuracy is computed based on these binary scores.
We test all three post-processing options in visual RAG and report the best result.
@@
TXT TAB CHA ALL TXT TAB CHA ALL
Textual RAG
LlamaParse-Premium 83.3 59.1 69.1 70.5 80.1 59.2 72.0 70.7
Qwen2.5-VL-72B 85.0 52.3 69.5 68.7 84.5 68.4 62.2 71.7
DocLayout-YOLO-Qwen2.5-VL-72B 61.7 25.0 28.6 38.4 (--30.3) 49.3 21.1 23.2 31.2 (--40.5)
Beehive-Qwen2.5-VL-72B 70.0 29.6 21.4 40.3 (--28.4) 61.3 40.8 26.8 43.0 (--28.7)
SCAN-Qwen2.5-VL-72B (ours) 81.0 70.5 73.8 75.3 ($+$6.6) 86.6 79.0 68.3 78.0 ($+$6.3)
Visual RAG
Qwen2.5-VL-7B 71.7 56.8 57.1 58.9 75.9 71.1 62.5 69.9
SCAN-Qwen2.5-VL-7B (ours) 75.0 63.6 57.1 65.3 ($+$6.4) 79.4 64.5 76.3 73.4 ($+$3.5)
Table 4: Textual and visual RAG results on Japanese datasets: BizMMRAG and Allganize.
@@
Qwen2.5-VL-7B 84.0 68.6 71.5 58.7 67.9 70.2
SCAN-Qwen2.5-VL-7B (ours) 86.1 68.6 73.2 61.7 87.8 75.5 ($+$5.3)
@@ 4.2 Textual RAG Evaluation Results
Our SCAN model can improve the performance of VLM-based text conversions. Even the strong baseline of Qwen2.5-VL-72B model achieves an impressive overall score of 31.1% (the ground truth is 36.1%), applying SCAN to Qwen2.5-VL-72B further improves the performance to 32.8%. The performance gains are larger when applying SCAN to OCR-specialized small VLMs. With GOT and Nougat, our SCAN improvements are 3.5% and 9.0%, respectively, enabling these smaller models to achieve competitive performance comparable to much larger-sized VLMs. This finding has important implications for deployment scenarios with computational constraints, suggesting that our semantic layout analysis approach can help bridge the efficiency-performance gap. The generation results exhibit similar improvement patterns. The results also indicate that SCAN’s enhancements for structured content elements such as reading order (RO), tables (TAB) and formulas (FOR) become increasingly significant. This suggests that the semantic segmentation approach is particularly valuable for preserving the relationships between elements that have spatial dependencies.
@@
On the other hand, applying SCAN slightly degrades the retrieval performance for some models. This is because retrieval represents a relatively simpler task within the RAG pipeline, primarily requiring the correct identification of keywords rather than comprehensive document understanding. In contrast, the subsequent question-answering stage demands precise and complete conversion of document content into text, where SCAN’s semantic layout analysis proves particularly advantageous.
@@
We can also see that conventional fine-grained document analysis methods substantially degrade overall performance, with DocLayout-YOLO Zhao et al. (2024b) reducing retrieval performance by 23.3% and overall performance by 8.3%. Beehive Auer et al. (2024) has the same trends that reduce 5.8% overall. These degradations are particularly severe for structured content types such as tables and charts. This proves that conventional layout analysis methods typically segment documents into small atomic regions, which also break the structure of documents. In contrast, our semantic box approach preserves the integrity of semantically coherent regions, maintains their holistic structure while still providing the organizational benefits of layout analysis. This preservation of semantic unity enables VLMs to process each region with full contextual awareness, leading to more accurate text conversion and, ultimately, superior RAG performance.
@@
BizMMRAG and Allganize. The upper section of Table 4 presents the textual RAG performance for Japanese document datasets. The results have the same trends as the English OHR-Bench evaluation, demonstrating that our SCAN methodology yields substantial improvements over general VLMs. Specifically, on the BizMMRAG dataset, our SCAN-enhanced approach demonstrates notable 6.6% improvements compared to the baseline Qwen2.5-VL-72B model: text accuracy decreased by 4.0%, while table increased by 18.2% and chart increased by 4.3%. We observe similar trends for the Allganize dataset. It is particularly noteworthy that our SCAN-enhanced method outperforms LlamaParse-Premium LlamaParse.AI (2025), a commercial product offered by LlamaIndex, specifically designed for document conversion. This comparison against a commercial solution validates the practical utility and competitive advantage of our research contribution. In addition, the result shows the same trends: using fine-grained document analysis, DocLayout-YOLO, and Beehive heavily degrades overall performance.
OHR-Bench BizMMRAG Allganize
semantic_box-only 75.1 65.3 72.1
concatenation 73.4 65.3 73.4
all-box 75.5 58.9 71.7
Table 6: Ablation study to compare the different post-processing strategies in visual RAG.
Setting # Input tokens # Output tokens Time (s)
Single-page 11,295 (prompt=255) 1,005 71.3
Multi-chunks 8,079 (prompt=1,071) 1,092 72.7
Table 7: Token numbers, processing time comparison of VLM text conversion.
@@ 4.3 Visual RAG Evaluation Results
OHR-Bench. Table 5 presents the result of OHR-Bench in visual RAG. When applying our SCAN approach to divide original pages into semantic chunks and performing visual RAG on these chunks, we observed an overall improvement of 5.3% compared to processing entire pages. While the Table task has the same performance as the baseline, we can see modest improvements in other categories, along with a significant enhancement in the Reading Order task. Recall that Reading Order problems require examining different paragraphs and articles to summarize answers. This demonstrates that dividing a page image into independent semantic chunks can reduce noise in reading order-related problems, enabling the model to provide more accurate responses.
BizMMRAG and Allganize. The lower section of Table 4 presents the visual RAG results for Japanese document datasets. The findings demonstrate that on both the BizMMRAG and Allganize benchmarks, our SCAN methodology exhibits substantial accuracy improvements. Specifically, SCAN improves 6.4% on BizMMRAG and 3.5% on Allganize. This result also shows that our SCAN approach enables the VLM to achieve significantly enhanced performance in Japanese VQA.
4.4 Ablation study of post-processing
Table 6 shows visual RAG results with different post-processing options. all-box performed the on OHR-Bench, while for BizMMRAG and Allganize, concatenation yielded better scores. Upon examining the OHR-Bench data, we found that the questions are generally unrelated to titles, headers, and other global boxes. This could explain why the all-box approach (which indexes different elements separately) and the semantic_box-only approach (which directly filters out title elements) performed relatively well. Conversely, we observed that queries in BizMMRAG and Allganize typically required consideration of both global information and main content, most likely causing the superior performance of the concatenation method for these datasets. For practical applications, we recommend experimenting with different approaches to select the best post-processing option.

### v2

@@ 3.3 Post-processing for RAG
We use the outputs of SCAN as a preprocessing step for both textual and visual RAG pipelines (Figure 2). Given a page image, SCAN predicts a set of semantic and global boxes, and we crop the corresponding sub-images from the original page. For textual RAG, we convert each sub-image into text using a VLM-based OCR model. We then concatenate the resulting texts into a single page-level sequence according to a reading order estimated from the box coordinates. In our implementation, we use a simple rule-based reading order prediction: boxes are sorted by the $y$-coordinate of the upper-left corner and then by the $x$-coordinate. The concatenated text is indexed and used as the document representation in downstream textual RAG systems. For visual RAG, we directly use all predicted boxes (semantic and global) as individual image chunks.
@@ 4 Experiments
4.1 Datasets and Settings
We use three datasets to evaluate the RAG performance: one English dataset, OHR-Bench ohr, and two Japanese datasets, Allganize allganize and our in-house BizMMRAG. Each dataset is used for both textual and visual RAG evaluation, which involves answering questions based on retrieved document content. As far as we know, OHR-Bench is the first benchmark to systematically evaluate the cascading impact of OCR on RAG systems. It enables step-by-step evaluation across OCR, retrieval, generation, and overall. Table 4 gives a detailed summary of these datasets. Section 4.1.1, Section 4.1.2, and Appendix E provide more details on the experimental settings.
Model Architecture Training Data
DiT DiT (304M) IIT-CDIP (42M) + PubLayNet (360K)
DocLayout-YOLO YOLO11-X (57M) DocLayNet (80k)
Beehive RT-DETR-L (43M) DocLayNet (80k)
SCAN RT-DETR-X (86M) Our Dataset (22k)
Table 5: Comparison of layout analysis models.
4.1.1 Textual RAG Setting
When evaluating English textual RAG performance, we use OHR-Bench ohr and follow the same evaluation protocol. We use BGE-m3 bgem3 and BM25 as retrieval models, and meta-llama/Llama-3.1-8B-Instruct and Qwen/Qwen2-7B-Instruct as answer models to generate answers according to the retrieved top-2 results. We use three metrics: (a) retrieval, which calculates LCS (Longest Common Subsequence) to measure evidence inclusion in retrieved content; (b) generation, which measures the F1-score of QA when provided with the ground truth page; and (c) overall, which calculates the F1-score of QA for the end-to-end RAG pipeline. The F1-score is calculated using precision and recall of common tokens between the generated result and the ground truth. The final scores will be the average score across the combination of two retrieval models and two answer models.
For the Japanese datasets BizMMRAG and Allganize, we employ intfloat/multilingual-e5-large as a retrieval model. For each query, the top-5 retrieved results are fed to the answer generation model, which is GPT-4o (gpt-4o-2024-08-06). To evaluate the generated answers, we adopt the LLM-as-a-judge framework llmasajudge, using GPT-4o to assign an integer score from 1 to 5 to each generated answer. Answers receiving a score greater than 4 are considered correct (assigned a value of 1), while others are considered incorrect (assigned a value of 0). Final accuracy is computed based on these binary scores.
To demonstrate the effectiveness of our SCAN method, we compare several text conversion methods: (1) using VLMs directly for text conversion without any layout analysis; (2) using fine-grained layout analysis methods, including DiT li2022dit, DocLayout-YOLO doclayoutyolo, and Beehive docling, followed by a VLM for text conversion; and (3) our SCAN method followed by a VLM for text conversion. For the VLMs, we use OCR-specialized models, GOT got, Nougat nougat, and olmOCR olmocr, as well as a general VLM, Qwen/Qwen2.5-VL-72B Qwen2.5-VL. For the layout analysis models, we list their architectures and training data in Table 5. Note that our SCAN model’s training data is much smaller than the other two layout analysis models. For the settings (2) and (3), we apply the post-processing described in Section 3.3.
@@ 4.1.1 Textual RAG Setting
DiT-GOT-580M 67.5 50.3 47.7 35.4 4.6 51.9 (++6.5) 46.4 35.0 25.5 19.6 14.8 35.0 (++7.2) 41.9 26.8 21.3 15.4 10.4 29.6 (++5.0)
DocLayout-YOLO-GOT-580M 60.4 45.5 43.2 32.8 4.4 46.8 (++1.4) 41.9 31.4 24.5 22.2 18.3 32.7 (++4.9) 38.5 23.9 20.2 16.1 12.1 27.5 (++2.9)
Beehive-GOT-580M 65.2 49.5 48.7 39.0 4.7 51.2 (++5.8) 45.7 33.3 26.4 23.7 28.1 35.9 (++8.1) 41.4 25.0 21.5 16.6 17.0 29.6 (++5.0)
SCAN-GOT-580M (ours) 68.5 54.3 50.7 36.6 5.3 53.9 (++8.5) 46.2 37.9 27.5 20.8 24.6 36.9 (++9.1) 41.9 28.4 22.5 16.0 17.3 30.8 (++6.2)
@@
DiT-Qwen2.5-VL-72B 76.9 57.7 55.6 44.6 5.4 59.7 (++0.1) 48.7 41.8 29.7 26.6 24.3 39.8 (++2.3) 44.8 32.0 24.5 20.3 16.0 33.5 (++2.4)
DocLayout-YOLO-Qwen2.5-VL-72B 63.5 12.4 36.0 11.7 5.4 36.3 (-23.3) 41.2 10.2 19.0 7.2 18.2 24.4 (-13.1) 38.4 10.0 16.3 7.7 12.5 22.4 (-8.7)
Beehive-Qwen2.5-VL-72B 73.3 16.0 42.3 16.3 6.1 42.6 (-17.0) 46.6 11.3 21.3 10.2 27.2 28.2 (-9.3) 43 10.9 19.5 8.6 16.6 25.3 (-5.8)
SCAN-Qwen2.5-VL-72B (ours) 75.7 56.6 57.3 40.6 6.5 58.9 (-0.7) 48.4 43.3 31.9 27.6 26.7 40.8 (++3.3) 44.4 31.9 26.6 20.6 17.7 33.8 (++2.7)
Table 6: Textual RAG results on OHR-Bench. Comparison of various OCR methods across different evaluation metrics. TXT, TAB, FOR, CHA, RO, and ALL represent text, table, formula, chart, reading order, and their average, respectively. The RO category includes questions that require identifying the correct reading order to associate information from separate paragraphs. Ground Truth indicates the performance when using the ground truth page text for retrieval and generation. The bold values indicate the best performance in each category.
4.1.2 Visual RAG Setting
We also apply OHR-Bench, BizMMRAG, and Allganize to evaluate visual RAG performance. We use ColQwen2-v1.0 colpali as an image retrieval model with top-5 retrieval and Qwen/Qwen2.5-VL-7B as an answer model.
Similar to the textual RAG setting, we compare three chunking methods: (1) using single-page images as a visual segment for retrieval and generation; (2) using layout analysis methods, DiT, DocLayout-YOLO, and Beehive, to chunk page images into layout-based chunks; and (3) our SCAN method to chunk page images into semantic chunks.
@@ 4.1.2 Visual RAG Setting
TXT CHA TAB ALL TXT CHA TAB ALL
Qwen2.5-VL-72B 85.0 52.3 69.1 68.8 84.5 68.4 62.2 71.7
DiT-Qwen2.5-VL-72B 75.0 54.6 50.0 59.9 (-8.8) 90.1 67.1 62.2 73.2 (++1.5)
DocLayout-YOLO-Qwen2.5-VL-72B 61.7 25.0 28.6 38.4 (-30.3) 49.3 21.1 23.2 31.2 (-40.5)
Beehive-Qwen2.5-VL-72B 70.0 29.6 21.4 40.3 (-28.4) 61.3 40.8 26.8 43.0 (-28.7)
SCAN-Qwen2.5-VL-72B (ours) 81.7 72.7 73.8 76.1 (++7.4) 85.9 85.5 72.0 81.1 (++9.4)
Table 7: Textual RAG results on Japanese datasets: BizMMRAG and Allganize.
@@
No chunking 84.0 68.6 71.5 58.7 67.9 70.2
DiT 80.7 63.8 62.2 51.9 66.0 64.9 (-5.3)
DocLayout-YOLO 72.2 57.9 58.3 47.5 62.4 59.6 (-10.6)
Beehive 73.2 60.1 64.2 43.8 87.6 65.8 (-4.4)
SCAN (ours) 86.0 70.0 73.5 63.4 86.3 75.8 (++5.6)
@@
BizMMRAG Allganize
TXT CHA TAB ALL TXT CHA TAB ALL
No chunking 71.7 56.8 57.1 58.9 75.9 71.1 62.5 69.9
DiT 61.7 59.5 63.6 61.6 (++2.7) 81.8 68.8 64.1 71.6 (++1.7)
DocLayout-YOLO 55.0 54.8 61.4 57.0 (++1.9) 69.5 68.8 65.8 68.0 (-1.9)
Beehive 66.7 45.5 52.4 54.8 (-4.1) 66.0 52.6 60.0 59.5 (-10.4)
SCAN (ours) 75.0 61.4 71.4 69.3 (++10.4) 84.4 67.1 75.0 75.5 (++5.6)
Table 9: Visual RAG results on Japanese datasets: BizMMRAG and Allganize.
@@ 4.2 Textual RAG Evaluation Results
Our SCAN model can improve the performance of VLM-based text conversions. Even the strong baseline of Qwen2.5-VL-72B model achieves an impressive overall score of 31.1% (the ground truth is 36.1%), applying SCAN to Qwen2.5-VL-72B further improves the performance to 33.8%. The performance gains are larger when applying SCAN to OCR-specialized small VLMs. With GOT, our SCAN’s improvement is 6.2 points, enabling this smaller model to achieve competitive performance comparable to much larger VLMs. This finding has important implications for deployment scenarios with computational constraints, suggesting that our semantic layout analysis approach can help bridge the efficiency-performance gap. The generation results exhibit similar improvement patterns. The results also indicate that SCAN’s enhancements for structured content elements such as reading order (RO) and tables (TAB) become increasingly significant. This suggests that the semantic segmentation approach is particularly valuable for preserving the relationships between elements that have spatial dependencies.
@@
On the other hand, applying SCAN slightly degrades the retrieval performance when using Qwen2.5-VL-72B. This is because retrieval is a relatively simpler task within the RAG pipeline, primarily requiring the correct identification of keywords rather than comprehensive document understanding. In contrast, the subsequent question-answering stage demands precise and complete conversion of document content into text, where SCAN’s semantic layout analysis proves particularly advantageous.
@@
We can also see that fine-grained document analysis methods, DocLayout-YOLO doclayoutyolo and Beehive docling, substantially degrade overall performance when used with Qwen2.5-VL-72B. The degradations are particularly severe for structured content types such as tables and charts. These conventional layout analysis methods typically segment documents into small atomic regions, which also break the structure of documents. In contrast, DiT and our semantic box approach improve the strong baseline of Qwen2.5-VL-72B. Although DiT is generally categorized as a conventional fine-grained document layout analysis method, its outputs include coarser segments compared to DocLayout-YOLO and Beehive, demonstrating that relatively coarser segments are more suitable for VLMs. Our SCAN further optimizes this level of granularity: it preserves the integrity of semantically coherent regions, maintains their holistic structure while still providing the organizational benefits of layout analysis. This preservation of semantic unity enables VLMs to process each region with full contextual awareness, resulting in more accurate text conversion and, ultimately, superior RAG performance.
It is notable that SCAN demonstrates high performance on English document benchmarks despite being trained exclusively on Japanese documents. This suggests that for common layout patterns, the impact of language may be less significant than the effectiveness of the coarse-grained segmentation approach itself.
@@
BizMMRAG and Allganize. Table 7 presents the textual RAG performance for Japanese document datasets. We include DiT, DocLayout-YOLO, and Beehive layout analysis models in our experiments, as the performance of layout analysis is not heavily dependent on language. The results have the same trends as the English OHR-Bench evaluation, demonstrating that our SCAN methodology yields substantial improvements over general VLMs. Specifically, on the BizMMRAG dataset, our SCAN-enhanced approach demonstrates notable 7.4 points improvements compared to the baseline Qwen2.5-VL-72B model: text accuracy decreased by 3.3 points, while chart accuracy increased by 20.4 points and table accuracy increased by 4.7 points. We observe similar trends for the Allganize dataset.
We also observe that the performance improvements are larger for Japanese datasets than for the English OHR-Bench. When using Qwen2.5-VL-72B, the gain was 2.7 points on OHR-Bench, but it increases to 7.4 points on BizMMRAG and 9.4 points on Allganize. In general, multilingual VLMs tend to demonstrate relatively higher performance in major languages such as English. Thus, while they can achieve reasonable performance on English documents even when the input images are complex, their performance is likely to degrade substantially on Japanese documents due to greater layout and linguistic complexity. This indicates that SCAN, which mitigates the input image complexity for VLMs, is especially effective for non-major languages such as Japanese, yielding even greater benefits than in English.
@@ 4.3 Visual RAG Evaluation Results
OHR-Bench. Table 8 presents the result of OHR-Bench in visual RAG. When applying our SCAN approach to divide original pages into semantic chunks and performing visual RAG on these chunks, we observed an overall improvement of 5.6 points compared to processing entire pages. We can see that SCAN approach is effective for every category. Especially, in the RO (reading order) task, our method achieves an impressive 18.4 points improvement. Recall that the RO task requires examining different paragraphs and articles to summarize answers. This demonstrates that dividing a page image into independent semantic chunks enables the system to retrieve only the relevant paragraphs, avoiding distractions from unrelated content on the same page.
BizMMRAG and Allganize. Table 9 presents the visual RAG results for Japanese document datasets. The findings demonstrate that on both the BizMMRAG and Allganize benchmarks, our SCAN methodology exhibits substantial accuracy improvements. Specifically, SCAN improves 10.4 points on BizMMRAG and 5.6 points on Allganize. This result also shows that our SCAN approach enables the VLM to achieve significantly enhanced performance in Japanese VQA.
Setting # Input Tokens # Output Tokens # Chunks Time (s)
No chunking 1,320.4 991.9 1.0 68.0
SCAN 9,683.1 2,515.0 12.4 56.3
Table 10: Processing cost and time comparison of VLM text conversion. We used 10 images randomly sampled from OHR-Bench, and the values in the table are averages over these 10 instances.